In Draco Corporation's first year of business, the following transactions affected its equity accounts.


- Issued 7,000 shares of $2 par value common stock for $48. It authorized 20.000 shares.

- Issued 1,750 shares of 12 %, $10 par value preferred stock for $53. It authorized 3,000 shares.

- Reacquired 350 shares of common stock for $60 each.

- Retained earnings is impacted by reported net income of $80,000 and cash dividends of $30,000.

Prepare the stockholders' equity section of Draco's balance sheet as of December 31. (Amounts to be deducted should be indicated by a minus sign.)
Answer #1
DRACO CORPORATION
Stockholders' Equity Section of the Balance Sheet
December 31
Preferred stock- $10 par value $17500
Paid in capital in excess of par- Preferred stock 75250
Common stock- $2 par value 14000
Paid in capital in excess of par- Common stock 322000
Retained earnings 50000
Less: Treasury stock -21000
Total stockholders' equity $457750

Preferred stock= 1750*$10= $17500

Paid in capital in excess of par- Preferred stock= ($53-10)*1750= $75250

Common stock= 7000*$2= $14000

Paid in capital in excess of par- Common stock= ($48-2)*7000= $322000

Retained earnings= $80000-30000= $50000

Treasury stock= 350*$60= $21000